# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v2

@@ Abstract
We build a dynamic multi-region model of climate and economy with emission permit trading among 12 aggregated regions in the world. We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments that were first outlined in the 2015 Paris Agreement and updated in subsequent years. Our model shows that the emission permit price reaches \$811 per ton of carbon by 2050. We demonstrate that a regional carbon tax is complementary to the global cap-and-trade system, and the optimal regional carbon tax is equal to the difference between the regional marginal abatement cost and the permit price.
@@ 1 Introduction
Climate change has been raising concerns of disastrous consequences, from rising sea levels to increases in the frequency and intensity of extreme weather events (Arias et al., 2021). International efforts to address climate change have evolved over the past three decades, commencing with the United Nations Framework Convention on Climate Change (UNFCCC) in 1992, followed by the Kyoto Protocol in 1997, the Paris Agreement in 2015, and the Glasgow Pact in 2021. The policy outlined in the Paris Agreement codified an objective of “limiting global average temperature rise to well below 2 degrees Celsius above pre-industrial levels, with a pursuit of limiting it to the 1.5 degrees Celsius” (Article 2). To accomplish this objective, 194 countries (parties) committed to the Nationally Determined Contributions (NDCs) specifying their emission mitigation goals, which need to be updated every five years by strengthening emission cuts with more intensive measures. Most recently, NDCs have been revisited in the Glasgow Climate Pact in November 2021, reaffirming the previous commitments made in the Paris Agreement and recognizing the need for more progressive efforts to attain the global target of 1.5 degrees Celsius.
@@
Various market-based approaches to reduce emissions have been proposed, among which the most well-known are carbon tax regimes and emission trading systems (ETSs). [^1] It has been documented that a total of 61 carbon pricing policies, consisting of 30 taxes and 31 ETSs, have been executed or are scheduled for future implementation around the world (Stavins, 2022). In a carbon tax regime, a tax is charged on firms for each unit of carbon emissions. Previous studies have shown that the optimal global carbon tax rate is determined at the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus, 2017). In an emission trading or a cap-and-trade system, on the other hand, the maximum amount of emissions is fixed in an economy, and agents can sell and purchase emission permits (allowances) at the market price. In the absence of transaction costs and uncertainty, properly designed tax and trading regimes are argued to yield equivalent outcomes in several key aspects (e.g., incentives for emission reductions, aggregate abatement costs, and carbon leakage) (Montgomery, 1972; Goulder and Schein, 2013; Stavins, 2022). In 2005, the European Union (EU) first adopted a legally binding emission trading system among the EU members, [^2] In addition to the EU ETS, other national or sub-national ETSs have been implemented or are in the process of development, including in Canada, China, Japan, New Zealand, South Korea, Switzerland, and the United States. For the information on most up-to-date policy in practice, see World Bank Carbon Pricing Dashboard (https://carbonpricingdashboard.worldbank.org/). and as of the summer of 2023, the permit price stands at about 90 euros per ton of Carbon Dioxide (CO2). [^3] Source of EU permit price: (https://tradingeconomics.com/commodity/carbon). Although a global-level ETS is currently not in place, there have been longstanding discussions on its implementation as an international regime to address climate change, as implied by Article 17 of the Kyoto Protocol and Article 6 of the Paris Agreement.
@@
In this study, we develop a dynamic multi-region model of climate and the economy under a hypothetical global ETS, and use it to project global prices of emission permits and the regional Social Cost of Carbon (SCC). In this integrated assessment framework, we formed 12 aggregated regions by grouping 198 countries around the world. Each region is allocated emission caps in accordance with the emission mitigation targets from the Paris Agreement and regional net zero commitments. Our main analysis involves solving a dynamic Nash equilibrium under noncooperation, where each region makes optimal dynamic decisions on emission abatement, emission permit purchases, and capital investment to maximize its social welfare given the optimal choices of other regions. Under the global ETS, regions with higher marginal abatement costs (MAC) are encouraged to purchase permits from the market to meet their emission cap constraints, whereas regions with lower MAC are incentivized to conduct more abatement and make profits by selling permits. The emission permit price is endogenously determined by annual supply and demand in the global permit market.
@@
Our study makes four contributions. The first contribution is demonstrating, both theoretically and numerically, that regional optimal taxes can be internally implemented under the global ETS. While carbon tax and ETS have often been considered as potentially interchangeable policies, we show that within the global ETS framework, it is optimal for the social planner of each region to introduce a carbon tax as a complementary policy tool to achieve its optimal level of net emissions. That is, a hybrid policy with both ETS and carbon tax is optimal. The necessity of a carbon tax arises from the discrepancy in optimal net emissions between the social planner’s choice and the market outcome under the ETS only (i.e., firms’ choice under no carbon tax). When optimal net emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional MAC and the global permit price under the noncooperative Nash equilibrium with the ETS, thereby ensuring firms’ choice under the tax is consistent with the social planner’s. We also numerically verify that the regional optimal carbon tax is equal to the regional SCC when net emissions are strictly positive.
@@
The second contribution of this study is a method to solve for the dynamic Nash equilibrium prices of emission permits under a global ETS. To the best of our knowledge, this is the first attempt to integrate an emission permit trading system with fully endogenous prices and emission abatements in a dynamic regional integrated assessment model (IAM) under noncooperation. Although some prior research has addressed regional integrated assessment models with a multi-region ETS, their solutions thus far have been obtained in a limited manner. One seminal work in IAMs is the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus and Yang, 1996; Nordhaus, 2010), but it did not provide a solution for the market-clearing equilibrium permit prices under noncooperation. Other studies have explored an ETS by extending the World Induced Technical Change Hybrid Model framework (WITCH; Bosetti et al., 2006), yet their permit prices and emission abatements are not fully endogenized as their prices are assumed to be equal to their regional MACs (De Cian and Tavoni, 2012; Massetti and Tavoni, 2012). While it has been long considered in the literature that the market equilibrium permit price is equalized to the regional MACs across all countries in an ETS regime (e.g., Keohane, 2009; Nordhaus, 2010), our study shows that this does not hold. Specifically, the assumption that the permit price equals the regional MACs ignores the inconsistency between forward-looking regional social planners’ and myopic firms’ choices under the ETS and no carbon tax. This can also be understood with an extreme case: if the emission caps are very large such that they are not binding for every region (which is equivalent to a case without emission caps), then both permit prices and trading quantities are zero, implying zero emission mitigation with the assumption. However, it is well known that even without emission caps, there are still nonzero emission mitigations and the optimal regional taxes are equal to the nonzero regional MACs under the Nash equilibrium (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
@@
We present a numerical algorithm for finding a Nash equilibrium solution where forward-looking regional social planners are making dynamic optimal decisions, and, importantly, show that the market clearing permit price is not equal to the regional MACs and the regional MACs are not equalized across regions. Under the baseline emission cap scenario based on the Glasgow Pact and the regional net zero commitments, our simulation results show that the oversupply of global permits in the initial decade results in zero permit prices, but by 2050, the emission permit price can reach up to \$811 per ton of carbon. Our results further show that regions such as the United States, India and the EU become major emission permit buyers, and other regions such as Russia and China become permit suppliers.
@@
The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Glasgow Pact, rendering our findings relevant for the real world scenario. We calibrate our model by fitting it to historical data and the recent predicted trends of regional climate damage, regional GDP, regional abatement costs, and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz, 2017), also known as the “Middle-of-the-Road” scenario. Besides this calibration, our model uses a stylized but stable climate system, called the Transient Climate Response to Emissions (TCRE; Matthews et al., 2009), which assumes that increases in the global averaged atmospheric temperature have a nearly linear functional dependence on cumulative carbon emissions. We show that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al., 2011). Due to its simplicity and effectiveness, the TCRE scheme has found applications in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas, 2017; Dietz and Venmans, 2019; Mattauch et al., 2020; Barnett et al., 2020. Moreover, Dietz et al. (2021) show that the TCRE scheme does not lead to a large difference in economic analysis with the seminal DICE framework (Nordhaus, 2017), compared to other more complicated climate systems.
@@
Our final contribution is a comprehensive comparative analysis of the roles played by cooperation, emission caps, and the global ETS. Our simulations reveal that global cooperation leads to significantly lower emissions and lower temperature rises. We also find that the global ETS leads to higher emissions under noncooperation, because regions with binding emission caps can purchase permits from regions with less restrictive caps, exploiting the excessively abundant emission permits in the first decade. This implies that it is important to maintain a stringent emission cap at the global level to make the ETS work efficiently. Moreover, we show that a stricter emission cap leads to a higher permit price and smaller regional SCCs for every region. For example, if all the regions achieve net zero emissions in 2050, the emission permit price can reach up to \$1,593 per ton of carbon by 2049. However, even with this strict emission restriction, the global temperature is still projected to rise by 1.61 degrees Celsius by the end of this century. These findings suggest that achieving the global target of limiting the temperature rise to 1.5 degrees Celsius requires even stronger emission commitments than those outlined in the Glasgow Pact. Our results are in qualitative agreement with the analysis of Rogelj et al. (2016), which found that the first NDC targets of the countries would result in a median warming of 2.1 to 3.1 degrees Celsius by 2100, and therefore stronger measures are needed to keep the warming under 1.5 degrees Celsius above the pre-industrial levels. Riahi et al. (2015) and Spencer et al. (2015) have similarly concluded that 2030 mitigation targets are insufficient to reach the greenhouse gas stabilization levels of 450 ppm.
We also find that under cooperation between the regions, the regional SCC is higher for a developed region compared to a developing one. That is, there are higher carbon taxes in developed regions under cooperation compared to developing regions because the marginal utility of consumption is higher in developing regions. The global aggregate welfare will be higher if developed regions can contribute more to mitigating emissions, thus reducing climate damages in developing regions. In the noncooperative setting, the regional SCCs of the developed and the developing regions are comparable.
@@
Our model framework is closely related to the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010), an extension of the global DICE model (Nordhaus, 2014) in a multi-region framework, capturing interactions between economic growth and climate systems. In Nordhaus (2010), carbon prices are obtained under a cooperative model where the global welfare is maximized with weights applied on regional utilities. Following the RICE model, a group of studies have explored climate policy in a dynamic multi-region framework under cooperation and noncooperation. For instance, Luderer et al. (2012) and Jakob et al. (2012) compare the long-term predictions of three regionalized energy-economy models, when certain environmental goals are set, like stabilizing the CO2 levels to 450 ppm. However, these works do not incorporate actual NDC and net-zero year targets of the countries to study the region-wise emission-economy pathways. van der Ploeg and de Zeeuw (2016) and Jaakkola and Van der Ploeg (2019) focus on stochastic components, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation. Hambel et al. (2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs, under certain model assumptions. Iverson and Karp (2021) study a Markov Perfect equilibrium in a dynamic game with a social planner deciding climate policies and non-constant discount rates. Cai et al. (2023) build a dynamic IAM for two economic regions (North and Tropic/South) to compute regional SCCs under cooperation and noncooperation. Nonetheless, none of these works examine the dynamics of an international ETS with an endogenous market price of emission permits under a noncooperative setting. Our study builds on the concept of regional SCCs defined in van der Ploeg and de Zeeuw (2016) and Cai et al. (2023), and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
@@
In the literature, ETSs have garnered attention as promising mechanisms for mitigating global emissions, along with carbon taxes, but there have been relatively limited attempts to comprehensively analyze ETSs as potential instruments for global climate policy. Among the most closely related studies are extended versions of the WITCH model (Bosetti et al., 2006): Massetti and Tavoni (2012) analyze an ETS for Asia in particular, where emission quotas are defined according to specific climate objectives, while De Cian and Tavoni (2012) study an ETS where countries can choose abatement and investment in innovation that lowers marginal abatement cost. As noted earlier, our solution mechanism differs from these studies and provides permit prices from a complete noncooperative dynamic optimization problem. Carbone et al. (2009) present another relevant study, constructing a computable general equilibrium model incorporating countries’ endogenous participation in an ETS and allocation of emission permits, solving for the equilibrium permit price. However, that study examines the ETS in a static setting, missing the critical dynamic aspects of climate change and lacking a connection with emission abatement decisions. Fischer and Springborn (2011) develop a dynamic stochastic general equilibrium model to compare outcomes of emission cap and tax policies, albeit without consideration of emission trading between regions. Other strand of studies have analyzed the potential efficiency gains from integrating regional ETSs into a global ETS (Habla and Winkler, 2018; Doda et al., 2019; Holtsmark and Weitzman, 2020; Holtsmark and Midttømme, 2021).
@@
More broadly, a substantial body of literature have examined ETSs from various perspectives. While in this study, we assume that emission permits are freely allocated following the voluntary commitments from the Paris Agreement, few studies have analyzed how the initial allocation of permits affects the equilibrium outcomes (Hahn and Stavins, 2011), or have compared allocation methods between auctioning and free allocation (Goulder and Parry, 2008; Goulder et al., 2010). Some have studied the relative efficiency of carbon policies under the existence of cost uncertainty, showing that carbon tax can be more efficient than ETS under certain conditions and vice versa (Weitzman, 1974; Stavins, 1996; Karp and Traeger, 2024). Some studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS (Hitzemann and Uhrig-Homburg, 2018; Fuss et al., 2018; Perino et al., 2022), China ETS (Goulder et al., 2022), or California ETS (Borenstein et al., 2019). An important issue with regional ETSs is carbon leakage, which refers to the shift of emission intensive production to regions outside the jurisdiction of the ETS. Previous studies have compared different policy instruments aimed at mitigating this problem and providing a level-playing field to the firms operating within the ETS (Ambec et al., 2024; Böhringer et al., 2014; Farrokhi and Lashkaripour, 2021; Fowlie and Reguant, 2022; Levinson, 2023).

### v3

@@ Abstract
We build a dynamic multi-region model of climate and economy with emission permit trading among 12 aggregated regions in the world. We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments that were first outlined in the 2015 Paris Agreement and later strengthened under the Glasgow Pact. Our model shows that the emission permit price reaches \$845 per ton of carbon by 2050, and global average temperature is expected to reach 1.7°C above the pre-industrial level by the end of this century. We demonstrate, both theoretically and numerically, that a regional carbon tax is complementary to the global cap-and-trade system, and the optimal regional carbon tax is equal to the difference between the regional marginal abatement cost and the permit price. We show that optimal regional carbon tax has significant heterogeneity between regions, and the tax needs to be implemented in both the developed and the developing regions.
@@ 1 Introduction
Climate change has been raising concerns of disastrous consequences, from rising sea levels to increase in the frequency and intensity of extreme weather events (Arias et al.,, 2021). International efforts to address climate change have evolved over the past three decades, commencing with the United Nations Framework Convention on Climate Change (UNFCCC) in 1992, followed by the Kyoto Protocol in 1997, the Paris Agreement in 2015, and the Glasgow Pact in 2021. The policy outlined in the Paris Agreement codified an objective of “limiting global average temperature rise to well below 2°C above pre-industrial levels, with a pursuit of limiting it to the 1.5°C” (Article 2). To accomplish this objective, 194 countries (regions) committed to the Nationally Determined Contributions (NDCs) specifying their emission mitigation goals, which need to be updated every five years by strengthening emission cuts with more intensive measures. Most recently, the Glasgow Climate Pact, adopted at the UN Climate Change Conference in Glasgow (COP26) in November 2021, revisited Nationally Determined Contributions (NDCs), reaffirming the previous commitments made in the Paris Agreement and recognizing the need for more stringent efforts to attain the global target of 1.5°C.
@@
Various market-based approaches to reduce emissions have been proposed, among which the most well-known are carbon tax regimes and emission trading systems (ETSs). [^1] It has been documented that a total of 61 carbon pricing policies, consisting of 30 taxes and 31 ETSs, have been executed or are scheduled for implementation globally (Stavins,, 2022). In a carbon tax regime, a tax is charged to firms for each unit of carbon emission. Previous studies have shown that the optimal global carbon tax rate is equal to the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus,, 2017). In an emission trading or a cap-and-trade system, on the other hand, the maximum amount of permissible emissions is fixed in an economy, and agents can sell and purchase emission permits (allowances) at the market price. In the absence of transaction costs and uncertainty, properly designed tax and trading regimes are argued to yield equivalent outcomes in several key aspects (e.g., incentives for emission reductions, aggregate abatement costs, and carbon leakage) (Montgomery,, 1972; Goulder and Schein,, 2013; Stavins,, 2022). In 2005, the European Union (EU) first adopted a legally binding emission trading system among the EU members, [^2] In addition to the EU ETS, other national or sub-national ETSs have been implemented or are in the process of development, including in Canada, China, Japan, New Zealand, South Korea, Switzerland, and the United States. For information on the most up-to-date policy in practice, see World Bank Carbon Pricing Dashboard (https://carbonpricingdashboard.worldbank.org/). and as of the winter of 2024, the permit price stands at about 70 euros per ton of carbon dioxide (CO2). [^3] Source of EU permit price: (https://tradingeconomics.com/commodity/carbon). Although a global-level ETS is currently not in place, there have been longstanding discussions on its implementation as an international regime to address climate change, as implied by Article 17 of the Kyoto Protocol and Article 6 of the Paris Agreement. Prior works have emphasized the need for a global ETS or linking existing regional ETSs (Mehling et al.,, 2018) and for expanding the coverage of carbon trading in specific industries to different regions so as to enhance the global mitigation efforts and reduce carbon leakages (Piris-Cabezas et al.,, 2018).
@@
In this study, we develop a dynamic multi-region model of climate and the economy under a global ETS with annual time steps, and use it to project global prices of emission permits and the regional SCC. In this integrated assessment framework, we form 12 aggregated regions by grouping 152 countries around the world. For these regions, we construct emission cap trajectories aligned with the latest NDC mitigation targets from the Paris Agreement and the Glasgow Climate Pact. These trajectories are based on near-term targets for 2025/2030 and long-term net-zero commitments for 2050–2070, and our model treats these trajectories as exogenous emission cap constraints. Assuming that all regions adhere strictly to their emission cap constraints without future revisions, [^4] In this paper, we assume that regions will not revise their NDC limits in future for the following reasons: (i) relaxing NDC limits could result in penalties from other regions or damage to the region’s reputation; (ii) reducing NDC limits would lead to a loss of benefits from selling emission permits or more cost from purchasing; and (iii) renegotiating NDC limits is often time-consuming and infrequent. our main analysis focuses on solving a dynamic Nash equilibrium under noncooperation, where each region makes optimal dynamic decisions on emission abatement, emission permit purchases, and capital investment to maximize its social welfare given the optimal choices of other regions. The emission permit price is endogenously determined by annual supply and demand in the global permit market.
@@
Our study makes four contributions. The first contribution is demonstrating, both theoretically and numerically, that regional optimal taxes can be internally implemented under the global ETS. While a carbon tax and ETS have often been considered as potentially interchangeable or substitutable policies, we show that within the global ETS framework, a hybrid policy combining both ETS and carbon tax is optimal: the social planner of each region needs to introduce a carbon tax as a complementary policy tool under the global ETS in order to achieve its optimal level of emissions. The necessity of a carbon tax arises from the discrepancy in optimal emissions between the social planner’s choice and the market outcome under an ETS-only regime (i.e., firms’ choice under the ETS with an exogenous emission cap but no carbon tax). Specifically, when optimal emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional marginal abatement cost (MAC) and the global permit price under the noncooperative Nash equilibrium with the ETS, thereby ensuring firms’ choice under the tax is consistent with that of the social planner’s. We also numerically verify that the regional optimal carbon tax is equal to the regional SCC provided emissions are strictly positive. Regional SCC is the present value of future regional climate damages resulting from additional global emissions.
@@
The second contribution of this study is a method to solve for the dynamic Nash equilibrium prices of emission permits under a global ETS. To the best of our knowledge, this is the first attempt to integrate an emission permit trading system with fully endogenous prices and emission abatements in a dynamic regional integrated assessment model (IAM) under noncooperation. Although some prior research has addressed regional IAMs with a multi-region ETS, their solutions thus far have been obtained in a limited manner. One seminal work in IAMs is the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus and Yang,, 1996; Nordhaus, 2010b, ), but it did not provide a solution for the market-clearing equilibrium permit prices under noncooperation. Other studies have explored an ETS by extending the World Induced Technical Change Hybrid Model framework (WITCH; Bosetti et al.,, 2006), yet their permit prices and emission abatements are not fully endogenized as their prices are assumed to be equal to their regional MACs (De Cian and Tavoni,, 2012; Massetti and Tavoni,, 2012). While it has been long considered in the literature that the market equilibrium permit price is equalized to the regional MACs across all countries in an ETS regime (e.g., Keohane,, 2009; Nordhaus, 2010b, ), our study shows that this does not hold. Specifically, the assumption that the permit price equals the regional MACs ignores the inconsistency between regional social planners’ choices internalizing the regional climate damages and firms’ choices ignoring the regional climate damages under the ETS regime. This can also be understood with an extreme case: if the emission caps are very large such that they are not binding for any region (which is equivalent to a case without emission caps), then both permit prices and trading quantities are zero, implying zero emission mitigation. However, it is well known that even without emission caps, there are still nonzero emission mitigations from the regional social planners’ perspectives and the optimal regional taxes are equal to the nonzero regional MACs under the Nash equilibrium (see, e.g., Nordhaus and Yang,, 1996; Cai et al.,, 2023). That is, regions’ social planners should impose an additional carbon tax on firms, contingent on their NDC commitments, such that the firms’ emission decisions are consistent with theirs.
@@
We present a numerical algorithm for finding a Nash equilibrium solution where forward-looking regional social planners are making dynamic optimal decisions, and importantly, we show that the market clearing permit price is not equal to the regional MACs and the regional MACs are not equalized across regions under the Nash equilibrium. Under the baseline emission cap scenario based on the Paris Agreement and Glasgow Pact commitments, our simulation results show that the oversupply of global permits in the initial years results in zero permit prices, but by 2050, the emission permit price can reach up to \$845 per ton of carbon. The global average temperature is expected to reach 1.7 degree Celcius above the pre-industrial level by the end of this century.
@@
The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Paris Agreement and the Glasgow Pact, making our findings directly relevant to real-world scenarios. We calibrate our model by fitting it to historical data and the recent predicted trends of regional climate damage, regional GDP, regional abatement costs, and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz,, 2017), also known as the “Middle-of-the-Road” scenario. Besides this calibration, our model uses a stylized but stable climate system, called the Transient Climate Response to Emissions (TCRE; Matthews et al.,, 2009), which assumes that increases in the global averaged atmospheric temperature have a nearly linear functional dependence on the cumulative carbon emissions. We show that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al.,, 2011). Due to its simplicity and effectiveness, the TCRE scheme has found applications in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas,, 2017; Dietz and Venmans,, 2019; Mattauch et al.,, 2020; Barnett et al.,, 2020. Moreover, Dietz et al., (2021) show that the TCRE scheme does not lead to a large difference in economic analysis with the seminal DICE framework (Nordhaus,, 2017), compared to other more complicated climate systems. Furthermore, our construction of emission cap pathways reflects the latest emission targets for 2030 and net-zero pledges for 2050-2070, contributing to the scant literature on evaluating the economic and environmental implications of these commitments (e.g., van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022).
@@
Our final contribution is a comprehensive comparative analysis of the roles of emission caps and the global ETS. We find that the global ETS leads to higher emissions under noncooperation, because regions with binding emission caps can purchase permits from regions with less restrictive caps, exploiting the excessively abundant emission permits in the initial years. This implies that it is important to maintain a stringent emission cap at the global level to make the ETS work efficiently. Moreover, we show that a stricter emission cap leads to a higher permit price for every region. For example, if all the regions achieve net zero emissions by 2050, the emission permit price can reach up to \$1,539 per ton of carbon by 2049. However, even with this strict emission restriction, the global temperature is still projected to rise by 1.62°C by the end of this century. These findings suggest that achieving the global target of limiting the temperature rise to 1.5°C requires even stronger emission commitments than those outlined in the Glasgow Pact. Our results are in qualitative agreement with the analysis of other IAMs (van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022), which find that the strengthened post-Glasgow NDC targets, encompassing both near-term and long-term goals, are insufficient to limit warming to 1.5°C above pre-industrial levels, but, if fully implemented, could limit warming to below 2°C. To ensure that our results are robust, we have performed sensitivity analysis by choosing alternative climate damage and abatement cost parameters obtained by fitting these parameters to other literature data.
@@
Our model framework is closely related to the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ), an extension of the global DICE model (Nordhaus,, 2014) in a multi-region framework, capturing interactions between economic growth and climate systems. In Nordhaus, 2010b , carbon prices are obtained under a cooperative model where the global welfare is maximized with weights applied on regional utilities. Following the RICE model, a group of studies have explored climate policy in a dynamic multi-region framework under cooperation and noncooperation. For instance, Luderer et al., (2012) and Jakob et al., (2012) compare the long-term predictions of three-region energy-economy models, when certain environmental goals are set, like stabilizing the CO2 levels to 450 ppm. However, these works do not incorporate actual NDC and net-zero year targets of the countries to study the region-wise emission-economy pathways. van der Ploeg and de Zeeuw, (2016) and Jaakkola and Van der Ploeg, (2019) focus on stochastic components, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation. Hambel et al., (2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs, under certain model assumptions. Iverson and Karp, (2021) study a Markov perfect equilibrium in a dynamic game with a social planner deciding climate policies and non-constant discount rates. Cai et al., (2023) build a dynamic IAM for two economic regions (North and Tropic/South) to compute regional SCCs under cooperation and noncooperation. Nonetheless, none of these works examine the dynamics of an international ETS with an endogenous market price of emission permits under a noncooperative setting. Our study builds on the concept of regional SCCs defined in van der Ploeg and de Zeeuw, (2016) and Cai et al., (2023), and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
@@
In the literature, ETSs have garnered attention as promising mechanisms for mitigating global emissions, along with carbon taxes, but there have been relatively limited attempts to comprehensively analyze ETSs as potential instruments for global climate policy. Among the most closely related studies are extended versions of the WITCH model (Bosetti et al.,, 2006): Massetti and Tavoni, (2012) analyze an ETS for Asia in particular, where emission quotas are defined according to specific climate objectives, while De Cian and Tavoni, (2012) study an ETS where countries can choose abatement and investment in innovation that lowers marginal abatement cost. As noted earlier, our solution mechanism differs from these studies and provides permit prices from a complete noncooperative dynamic optimization problem. Carbone et al., (2009) present another relevant study, constructing a computable general equilibrium model incorporating countries’ endogenous participation in an ETS and allocation of emission permits, solving for the equilibrium permit price. However, that study examines the ETS in a static setting, missing the critical dynamic aspects of climate change and lacking a connection with emission abatement decisions. Fischer and Springborn, (2011) develop a dynamic stochastic general equilibrium model to compare outcomes of emission cap and tax policies, albeit without consideration of emission trading between regions. Other strand of studies have analyzed the potential efficiency gains from integrating regional ETSs into a global ETS (Habla and Winkler,, 2018; Doda et al.,, 2019; Holtsmark and Weitzman,, 2020; Holtsmark and Midttømme,, 2021; Mehling et al.,, 2018).
@@
More broadly, a substantial body of literature has examined ETSs from various perspectives. While in this study, we assume that emission permits are freely allocated to the regions following the voluntary commitments from the NDCs, several studies have analyzed how the initial allocation of permits affects the equilibrium outcomes (Hahn and Stavins,, 2011), or have compared allocation methods between auctioning and free allocation (Goulder and Parry,, 2008; Goulder et al.,, 2010). Other works have studied the relative efficiency of carbon policies under cost uncertainty, showing that carbon tax can be more efficient than ETS under certain conditions, and vice versa (Weitzman,, 1974; Stavins,, 1996; Karp and Traeger,, 2024). In another strand of literature, a number of studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS (Hitzemann and Uhrig-Homburg,, 2018; Fuss et al.,, 2018; Perino et al.,, 2022), China ETS (Goulder et al.,, 2022), or California ETS (Borenstein et al.,, 2019). An important issue with regional ETSs is carbon leakage, which refers to the shift of emission intensive production to regions outside the jurisdiction of the ETS. Previous studies have compared different policy instruments aimed at mitigating this problem and providing a level-playing field to the firms operating within the ETS (Ambec et al.,, 2024; Böhringer et al.,, 2014; Farrokhi and Lashkaripour,, 2021; Fowlie and Reguant,, 2022; Levinson,, 2023).